OpenStudy (anonymous):

would you know if this linear equation intersects or a parallel line? if it intersects, where does it intersects? 2x-y=2 and y=3x-4

OpenStudy (anonymous):

There are two line rewritten 1) 2x-y=2 2) 3x-y=4

OpenStudy (anonymous):

slopes are different so they intersect

OpenStudy (anonymous):

One is not a factor of another so we know that they are not parellel

OpenStudy (anonymous):

first one is \[y=2x-2\] slope is 2 second line is \[y=3x-4\] slope is 3

OpenStudy (anonymous):

To find point of intersection set them equal to each other 2x-2=3x-4 and solve for x

OpenStudy (anonymous):

so yes, they intersect. if you want to know where, set them equal and solve via \[2x-2=3x-4\] \[-2+4=3x-2x\] \[-2=x\]

OpenStudy (anonymous):

what imranmeah said

OpenStudy (anonymous):

thank you...can i still ask ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

why is that -2=x...? is it not 2=x?

OpenStudy (anonymous):

Actually you are right; it is 2

OpenStudy (anonymous):

can i ask still?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

you guys answered this so fast...mind teaching me what is the shortest method?..because i tried to answer ...it took me almost 3 hours but did not answer one...so, can u help ,e?

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

i still cant understand and i want to understand

OpenStudy (anonymous):

If you are given two line and asked if they intersect,first thing you need to check is if they are parellel.

OpenStudy (anonymous):

Because lines always intersect each other unless they are parellel

OpenStudy (anonymous):

how to check if they are parallel or not?

OpenStudy (anonymous):

Parellel lines have same slope, so following lines are parellel y=3x+2 y=3x-5 y=3x+10 See same slope

OpenStudy (anonymous):

yes clear...with that

OpenStudy (anonymous):

with what I just said

OpenStudy (anonymous):

but i was asked actually to solve those systems by graphing...

OpenStudy (anonymous):

can u help me find the ordered pairs for me to graph?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

Just graph two line and see where they cross

OpenStudy (anonymous):

but i dont have given pairs...i need to know the pairs ,right?

OpenStudy (anonymous):

what pairs?

OpenStudy (anonymous):

of the x and y

OpenStudy (anonymous):

example : 2x-y=2 y=3x-4

OpenStudy (anonymous):

1)y=2x-2 2)y=3x-4 To graph 1) I would graph start at y intercept of -2 and move up with slope of 2

OpenStudy (anonymous):

where do they intersect?....

OpenStudy (anonymous):

ok i am following...and?

OpenStudy (anonymous):

Do the same with second one. After you have graph two line, see where they cross; that's the intersetion point

OpenStudy (anonymous):

why move up with a slope of u? meaning, in the y axis?

OpenStudy (anonymous):

i have a question....i have another example: x-2y=6 and y=x+1

OpenStudy (anonymous):

in the book it will intersect at one point whose coordinates are -8 and -7...with the first line x-2y=6 it has a pair of 0 and -3 and 6 and 0....the next line y=x+1 it has pairs of -1 and 0 and 0 and 1....where did those pairs come from?

OpenStudy (anonymous):

that is where i dont undestand

OpenStudy (anonymous):

If you plug in 0 for x you should get -3 for y x-2y=6 0-2y=6 y=-3

OpenStudy (anonymous):

Plug in 6 for x-2y=6 in place of x and you get get y=0 6-2y=6 -2y=0 y=0

OpenStudy (anonymous):

ya, but how did they get that? it is not given in the book

OpenStudy (anonymous):

You choose a x value of your liking and figure out what y will be

OpenStudy (anonymous):

oh..that is not a problem for me to solve?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

wow....how come the teacher did not say that?

OpenStudy (anonymous):

one more? what about the slope?

OpenStudy (anonymous):

what about it?

OpenStudy (anonymous):

how would i know if the slope is positive or negative?

OpenStudy (anonymous):

y=3x+2 positive slope y=-3x+2 negative slope you look that the number next to x

OpenStudy (anonymous):

nex to x...that is 2?

OpenStudy (anonymous):

Sorry "number with x"

OpenStudy (anonymous):

ok...will look at first the samples and will ge tback to u if i am in trouble

OpenStudy (anonymous):

by the way, where are u from? are u still going to school? u are such a smart boy...thank u